### v1

# Neuromorphic Computing with Microfluidic Memristors

Abstract
Conical microfluidic channels filled with electrolytes exhibit volatile memristive behavior, offering a promising foundation for energy-efficient, neuromorphic computing. Here, we integrate these iontronic channels as additional nonlinear element in nonlinear Shinriki-inspired oscillators and demonstrate that they exhibit alternating chaotic and non-chaotic dynamics across a broad frequency range. Exploiting this behavior, we construct XOR and NAND gates by coupling three “Memriki” oscillators, and we further realize the full set of standard logic gates through combinations of NAND gates. Our results establish a new paradigm for iontronic computing and open avenues for scalable, low-power logical operations in microfluidic and bio-inspired systems.
I Introduction
The era we live in has been proposed the Silicon age [1], defined by semiconductor transistors that have revolutionized data processing, accelerating it by several orders of magnitude. This technological advancement has enabled groundbreaking inventions, including computers, the internet, and, more recently, artificial intelligence (AI). However, advancements in AI are rapidly approaching a critical power bottleneck [2]. A key factor driving this challenge is the high energy cost caused by Ohmic losses of frequent data transfer between physically separated processing and storage units in conventional computing architectures, a limitation known as the von Neumann bottleneck [3]. AI power consumption is currently doubling every four to six months [4], which would surpass global energy production by the 2030s. This trend highlights the pressing need for novel computing paradigms. To enhance energy efficiency and meet the increasing demands for computing power in the era of AI and big data, a fundamentally different approach is required.
Inspired by the most energy-efficient computer known – the biological brain – the field of neuromorphic computing emerged [5]. The human brain, an intricate biological neural network comprising approximately 1011 neurons and 1015 synapses, operates with a remarkable efficiency at just $20\,\mathrm{W}$ of power. A key subfield of neuromorphic computing focuses on developing devices that mimic the computational functions of the brain. One such device is the memristor, first theoretically proposed by Chua in 1971 as the fourth basic circuit element [6]. Memristors act as memory elements, with a resistance that depends on past voltage or current, allowing them to retain information. In 2008, Strukov et al. experimentally demonstrated a solid-state memristor, paving the way for its use as an artificial synapse in neuromorphic systems [7], with present-day applications in for instance crossbar arrays [8, 9]. However, silicon-based technologies typically rely on a single type of information carrier: electrons (or their absence (holes)). In contrast, the biological brain utilizes ions dissolved in water as information carriers, allowing for a diverse range of soluble information carriers, such as various species of ions and small molecules. A key distinction between solid-state electronics and ion-mediated biological systems lies in their energy efficiency, with biological systems operating orders of magnitude more efficiently, albeit at much lower clock frequencies. This contrast has led to the emergence of iontronics, a subfield of neuromorphic computing focused on developing electronic circuit elements where ions dissolved in water serve as information carriers [10]. Just as solid-state memristors have been widely explored, researchers have pursued iontronic analogs [11]. Memristive effects in charged conical nanopores were first observed in 2010 and 2012 [12, 13], paving the way for the development of diverse nanofluidic memristors based on various mechanisms. These include electric double layer polarization [13, 14, 15, 16], salinity gradients yielding negative differential resistance behavior [17], electro-wetting [18, 19], structural and conformational changes of nanopores [20, 21], the ionic analogue [22] of electronic Coulomb blockade [23], mixtures of and interfaces between ionic liquids and water [24, 25], and specific polyelectrolyte-ion interactions [26].
The next step towards neuromorphic computing is integrating iontronic devices into electric circuits. Key questions include: Can these networks emulate neural functions, perform Boolean logic operations, or execute simple machine learning tasks [27]? While the use of iontronic elements in large-scale networks is still in its infancy, initial progress has been made in implementing basic logic gates such as AND, OR [28], NAND and NOR [29], and material implication IMP [15] gates. However, a complete set of logic gates based on (memristive) iontronics has yet to be realized.
In this work, we bridge iontronics and memristive computing by constructing logic gates using microfluidic memristors [30]. In addition to microfluidic memristors, we base our logic gates on Shinriki-inspired oscillators [31]. Shinriki oscillators are simple electric circuits, first studied by Shinriki et al., that can show chaotic behavior depending on the exact system parameters. As the human brain operates at the edge of chaos [32, 33], we designed our circuits to inherit this property.
We first provide a brief introduction to volatile microfluidic memristors as descibeed in Ref. [30], and then explore their integration into electric circuits that form Shinriki-inspired (non-linear) oscillators [31]. Finally, we demonstrate how these oscillators can be combined to realize the full set of logic gates, thereby advancing the use of iontronics for neuromorphic and Boolean logic applications.
II Microfluidic iontronic memristor
Figure 1: (a) Schematic of a conical channel filled with an electrolyte (cations in green, anions in red) with a length $L$, base radius $R_{b}$, tip radius $R_{t}$, bulk ion densities $\rho_{\pm}$, and surface charge $e\sigma$. A time-dependent voltage $U(t)$ is applied across the channel. (b) Limit cycles of the conductance-voltage relation for three different driving frequencies $f=0.4\,\mathrm{Hz}$ (blue), $f=40\,\mathrm{Hz}$ (orange), and $f=4000\,\mathrm{Hz}$ (green) of a sinusoidal potential with an amplitude of 1 Volt. (c) Limit cycles of the current-voltage relation for the same frequencies and parameters as in (b).
In this study, we focus on volatile microfluidic memristors as presented in Ref. [30] and illustrated in Fig. 1(a). This conical channel, with a length $L$, has a wide opening at $x=0$ and a narrow tip at $x=L$, and connects two reservoirs of an aqueous 1:1 electrolyte at a total bulk ion concentration $2\rho_{b}$ at room temperature. The channel can carry an ionic electric current when driven by an applied potential difference between the two reservoirs. The combination of the conical channel geometry and a negative surface charge $\sigma$ on the channel walls results in current rectification when a static voltage drop $U$ is applied [34]. This phenomenon was theoretically explained in terms of a nontrivial $U$-dependent steady-state salt concentration profile $\rho_{s}(x;U)$ for $x\in[0,L]$ that build up within the channel [35]. This concentration profile is caused by the much stronger electric field near the tip compared to the base and leads to ionic depletion or accumulation depending on the polarity of the potential. Consequently, the steady-state electric conductance $g_{\infty}(U)$ depends non-trivially on $U$ and exhibits a hysteresis loop in the current-voltage relation when a time-dependent oscillating potential $U(t)$ is applied with a period on the order of the build-up time $\tau$ of the ionic concentration profile [30]. The time-dependent conductance $g(t)$ of such a memristive channel was calculated from numerical solutions of the Poisson-Nernst-Planck-Stokes (partial differential) equations for the coupled transport of charge, water, and ions [30]. In a large parameter regime $g(t)$ was also found to be accurately approximated by the following (ordinary differential) equation of motion 111this is only an approximation of the steady-state conductance, but Kamsma et al.  found that it is reasonable within the applied voltage range [30, 44]
$\tau\dot{g}(t)=g_{\infty}(U(t))-g(t)\mbox{ with}$ (1)
$g_{\infty}(U)=g_{0}\int_{0}^{L}\frac{\rho_{s}(x,U)}{2\rho_{b}L}\mathrm{d}x,$ (2)
where the functional form of the static conductance $g_{\infty}(U)$ is determined by the channel geometry and the electrolyte properties, that we keep fixed here however such that its numerical evaluation for different $U$ is straightforward. The parameter $g_{0}$ represents the zero-field conductance, which we use as the unit of conductance below. Throughout this work we will use the (physically realistic) set of system parameters of Ref. [30], which are all given, together with detailed expressions for $g_{\infty}(U)$, in Section VII. Of particular importance is the characteristic time scale $\tau\approx 5\,\mathrm{ms}$ of our parameter set, which is the memory retention time of the microfluidic memristor. We note, however, that the dynamic response of the conical channel is qualitatively independent of the details of the system parameters.
We consider a sinusoidal voltage $U(t)=U_{0}\sin(2\pi ft)$ with fixed amplitude $U_{0}=1\,\mathrm{V}$ and several frequencies $f$ applied over the microfluidic channel to analyze its frequency-dependent dynamic response resulting from numerical solutions of Eq. (1). Focusing on the limit cycle in which all transients have decayed, we present in Fig. 1 parametric plots for three different frequencies, $(U(t),g(t))$ in (b) and $(U(t),I(t))$ in (c), where the current is defined as $I(t)=g(t)U(t)$. Depending on the driving frequency, the electric response of the channel falls into one of three regimes: (i) At low frequencies $f\ll 1/\tau$, exemplified here by $f=0.4\,\mathrm{Hz}$ (blue), the conductance $g(U(t))$ closely follows the steady-state function $g_{\infty}(U(t))$ in (b), as the driving voltage changes slowly enough for the concentration profile to fully build up at all times. In this low-frequency regime, the electric current $I(t)$ in (c) shows strong rectification, approaching nearly full diodic behavior, with significantly higher current at negative voltages than at positive ones. (ii) At high frequencies $f\gg 1/\tau$, exemplified here by $f=400\,\mathrm{Hz}$ (green), the sinusoidal voltage oscillates too rapidly for the ionic concentration profile to develop, leaving it essentially spatially and temporally constant. As a result, the conductance is essentially Ohmic with constant $g(t)\approx 1.07\,\mathrm{g_{0}}$ in (b) and $I(t)$ linear in $U(t)$ in (c). One might have expected the conductance in this high-frequency regime to be exactly equal to $g_{0}$. However, the nonlinearity of $g_{\infty}(U)$ results in a nonzero difference between the static conductance at the time-averaged voltage $g_{0}$ and the time-averaged steady-state conductance for the time-dependent voltage, yielding $\langle g_{\infty}(U(t)\rangle\approx 1.07\,\mathrm{g_{0}}$. (iii) At intermediate frequencies $f\sim 1/\tau$, exemplified here by $f=40\,\mathrm{Hz}$ (orange), the driving voltage varies on a timescale comparable to the memory timescale of the channel. As a result, hysteresis emerges in both the conductance (b) and the current (c) since the ionic concentration profile can partially build up and break down but not fully equilibrate. This leads to a lower channel conductance when the potential was recently positive and a higher conductance when it was recently negative.
In the remainder of this work, we measure all quantities relative to the steady-state conductance $g_{0}$ of the memristor at $0\,\mathrm{V}$. We use Volts ($\,\mathrm{V}$) and seconds ($\,\mathrm{s}$) as our units for potential and time, respectively. Specifically, resistance is expressed in $\,\mathrm{g_{0}}^{-1}$, current in $\,\mathrm{g_{0}V}$, capacitance in $\,\mathrm{g_{0}s}$, and inductance in $\,\mathrm{s}/\mathrm{g_{0}}$. We solve the equations of motion for all circuits in this paper using the circuit simulation software ACME.jl based on [37]. For all simulations, we use a time step $\delta t=10^{-5}\,\mathrm{s}$, except for the calculations in Fig. 1 (b) and (c), where we use $\delta t=10^{-5}/f$, to ensure sufficient resolution for the high frequency cases. The code for reproducing the simulations is provided in Ref. [38].
III Characterization of a Single Memriki oscillator
To leverage the volatile behavior of the memristors of Section II, we incorporate two of them into a nonlinear electrical circuit with an intrinsic oscillatory timescale similar to that of the memristors. The circuit design, shown in Fig. 2, is heavily inspired by the well-known Shinriki-oscillator [31], albeit with a slightly modified topology and with iontronic memristors $M_{1}$ and $M_{2}$ rather than diodes. We propose the name Memriki oscillator for this circuit. The circuit, driven by a sinusoidal source potential $U_{in}$ with frequency $f$, consists of a negative differential resistance $\mathrm{NDR}$, two anti-parallel memristors in series with a (weakly damped) LRC oscillator with eigen frequency $1/(2\pi\sqrt{L_{o}C_{o}})$, and a resistor $R_{1}$ and capacitor $C_{1}$ that are both parallel to the memristor-oscillator part. Based on the original work of Shinriki et al.  [31], the $\mathrm{NDR}$ is an active element with a characteristic function given by $R(U)\,\mathrm{g_{0}}=(U/\,\mathrm{V})^{3}-3U/\,\mathrm{V}$. This implies that it behaves as an active amplifier rather than a passive resistor within the ±$\pm 1\,\mathrm{V}$ voltage range around $0\,\mathrm{V}$. The parallel resistor and capacitor, with $C_{1}=10^{-4}\,\mathrm{g_{0}s}$ and $g_{0}R_{1}$ of order unity, can serve as tunable parameters that significantly alter the dynamics of the circuit. The electric specifications for the oscillatory part of the circuit read $C_{o}=10^{-3}\,\mathrm{g_{0}s}$, $L_{o}=10^{-3}\,\mathrm{s}/\mathrm{g_{0}}$ and $R_{o}=10^{-3}\,\mathrm{g_{0}}^{-1}$, ensuring that $R_{o}\ll R_{1}$. As a result, the eigen period of the weakly damped oscillator is approximately $2\pi\sqrt{L_{o}C_{o}}\approx 6\,\mathrm{ms}$, which is comparable to the memory time of the two memristors, $\tau\approx 5\,\text{ms}$. While we describe and denote all circuit elements (except for the memristors) as conventional electrical components, we note that microfluidic counterparts exist for each of them [39, 40]. This enables a purely microfluidic realization of this Memriki oscillator.
The dynamic response of the Memriki oscillator varies dramatically depending on the precise values of the resistor $R_{1}$ and the driving frequency $f$. For instance, it can exhibit behaviors ranging from simple driven oscillations to chaos to subharmonic oscillations, as illustrated in  Fig. 3. A comprehensive study of the circuit’s full dynamics across a broader parameter space is beyond the scope of this paper. Instead, we focus here on the coupling of multiple Memriki oscillators.
Figure 2: Schematic of a single Memriki oscillator. The capacitor $C_{o}$, inductor $L_{o}$ and resistor $R_{o}$ together form a damped oscillator. The circuit is driven by the input voltage $U_{\mathrm{in}}$, measured relative to the reference voltage $U_{\mathrm{ref}}$. A negative differential resistance $\mathrm{NDR}$ introduces nonlinearity into the circuit. The output voltage $U_{\mathrm{out}}$ is measured, stabilized by $C_{1}$, and slowly relaxed towards $U_{\mathrm{ref}}$ by $R_{1}$. Two iontronic memristors $M_{1}$ and $M_{2}$ are added between $U_{\mathrm{out}}$ and the oscillatory section to introduce further nonlinearity.
Figure 3: Dynamics of the driven Memriki oscillator from Fig. 2 for three different combinations of resistor $R_{1}$ and driving frequency $f$, with all other circuit elements specified in the text. The top panels (a)-(c) show the time series of the input voltage $U_{\mathrm{in}}=0.25\sin(2\pi ft)\,\mathrm{V}$ (orange) and the output voltage $U_{\mathrm{out}}$ (blue) after 100 periods. The bottom panels (d)-(f) show parametric plots of $U_{\mathrm{out}}$ versus $U_{\mathrm{in}}$ over 1000 periods, revealing stable periodic orbits in (d) and (f) separated by a chaotic, non-periodic regime in (e). Thus, in (a,d) the Memriki oscillator follows the input signal and oscillates at the driving frequency $f$. In (b,e), the response exhibits chaotic behavior, while in (c,f), the response is subharmonic, oscillating at a frequency of $f/3$.
IV Coupling Memriki oscillators
The most fundamental computations involve combining two inputs to produce a single output. To achieve this, we couple three Memriki oscillators, represented by the three “boxes” in the circuit diagrams of Fig. 4 (a) and (c). Two oscillators serve as inputs, while the third acts as an output. In Fig. 4(a), we see that the outputs of the two input oscillators are capacitively connected to the input of the output oscillator, which prevents any DC current from flowing from the inputs to the output. When the circuit of Fig. 4(a) is driven by input voltages 00 and $1\,\,\mathrm{V}$, representing low and high logic levels, along with Gaussian noise with a standard deviation of $0.003\,\mathrm{V}$, the output of the circuit turns out to correspond to a logic XOR-like gate. For $R_{1}=0.4\,\mathrm{g_{0}}^{-1}$, this is shown in the voltage time trace of Fig. 4(b), with blue and orange representing the four consecutive input combinations (10, 00, 01, 11) and green the output. That is, the circuit oscillates when the two inputs differ and remains quiescent when they are the same. This XOR-type behavior can be interpreted as performing a nonlinear classification task, one of the most fundamental classification tasks in machine learning [41]. For reasons that will become clear in the next section we refer to this as an XOR-pre-gate.
The NAND-gate is the primary objective when building logic gates, as all other gates can be constructed from combinations of NAND. We realize a NAND-pre-gate by inserting two additional operational amplifiers into the circuit as shown in Fig. 4(c). These amplifiers function as impedance changers, ensuring sustained oscillations in the third oscillator when both inputs are low. The resulting output voltage, shown in green in Fig. 4(d) for the same four input combinations (blue and orange) as in (b), oscillates for the input states 10, 00, and 01, but remains quiescent for input 11. Therefore, we classify this circuit as a NAND-pre-gate.
Figure 4: Circuit diagram of a XOR-pre-gate (a) and a NAND-pre-gate (c) constructed by combining three Memriki oscillators of Fig. 2 (here represented by the three boxes), either only coupled capacitively (a) or with two additional operational amplifiers (c). Time traces of the two input voltages (blue and orange) representing the four consecutive binary states 10, 00, 01, 11. The resulting quasi-binary output voltage (green), which is either quiescent or oscillating, is shown in (b) for circuit (a), revealing an XOR behavior, and in (d) for circuit (c), exhibiting a NAND behavior. The corresponding truth tables are shown alongside the circuit diagrams.
To use the output of one logic gate as the input for another, the output must be compatible with the input requirements, in this case static signals of $0\,\mathrm{V}$ and $1\,\mathrm{V}$. To achieve a proper NAND gate, we must transform the oscillatory output of our NAND-pre-gate into a constant output of $1\,\mathrm{V}$ while its quiescent (0) output for the 11 input remains unaffected. This transformation is achieved by passing the output voltage of our NAND-pre-gate through the coupler shown in the circuit diagram of Fig. 5. In this coupler, the output voltage first passes through operational amplifier OA1, which strengthens the signal and prevents drawing any current from the input gates. The amplified signal is then passed through a full-wave bridge rectifier composed of four diodes, followed by time-averaging with a low-pass filter with time constant $R_{LP}C_{LP}=0.5\,\mathrm{ms}$. Finally, the signal is amplified by AO2 to achieve the desired $1\,\mathrm{V}$ logic level.
Figure 5: Coupler for converting oscillations into constant logic levels. The oscillatory input voltage difference between $U_{\mathrm{in}}^{+}$ and $U_{\mathrm{in}}^{-}$ is first amplified by the left operational amplifier, then rectified by a full-wave bridge rectifier composed of four diodes. The rectified signal is subsequently smoothed by a low-pass filter consisting of $R_{\mathrm{LP}}$ and $C_{\mathrm{LP}}$ and finally amplified to the desired logic level by the right operational amplifier. The operational amplifiers in this circuit are modeled by Eq. 3, with a maximum voltage of $U_{\mathrm{cc}}=1\,\mathrm{V}$.
The two operational amplifiers used in the coupler convert the input voltage $U_{\mathrm{in}}$ into an output voltage given by
$U_{\mathrm{out}}=\Delta U\tanh\left(\frac{GU_{\mathrm{in}}}{\Delta U}\right),$ (3)
where $G$ is the gain factor (for small inputs), set to $G=15$ for $\mathrm{OA}_{1}$ and to $G=35$ for $\mathrm{OA}_{2}$. The parameter $\Delta U=1.0\,\mathrm{V}$ defines the range of the amplifier, allowing amplification up to ±$\pm 1\,\mathrm{V}$. The low-pass RC filter consists of a resistance $R_{\mathrm{LP}}=10^{-2}\,\mathrm{g_{0}}^{-1}$ and a capacitance $C_{\mathrm{LP}}=5\times 10^{-2}\,\mathrm{g_{0}s}$, while the diodes are modeled with a saturation current of $0.3\,\mathrm{g_{0}V}$ and an emission factor $\eta=1$ [42].
With the addition of this coupler circuit, the NAND-pre-gate becomes a true NAND-gate.
V Combining gates
The NAND gate is well known to be functionally complete, which means that any other logic gate can be constructed from it. In particular, an OR gate can be constructed using three NAND-gates, as shown in Fig. 6(a). The Boolean logic for the OR gate follows $\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{1}}\cdot\overline{U_{\mathrm{in}}^{2}\cdot U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}}\cdot\overline{U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}}}+\overline{\overline{U_{\mathrm{in}}^{2}}}=U_{\mathrm{in}}^{1}+U_{\mathrm{in}}^{2}$, where $\cdot$ denotes the logic AND, $+$ the logic OR, and an overline indicates negation. In our circuits, each input signal $U_{\mathrm{in}}^{1}$ and $U_{\mathrm{in}}^{2}$ is passed through the combined gates twice to ensure independent noise sources for all our Memriki gates. If identical signals were used, the differential input to the final gate would cancel out, preventing the oscillations from self-exciting. To further aid self-excitation, we also add Gaussian noise with a standard deviation of $0.003\,\mathrm{V}$ to the inputs of the last gate.
The voltage time trace in Fig. 6(c) displays the output (green) of the OR gate for the same four consecutive input signals (blue and orange) used previously. This confirms that the gate functions correctly as a logic OR gate, processing weakly noisy static inputs. The response time of the gate is approximately $0.1\,\mathrm{s}$ for relatively slowly varying input signals.
As is well known, the circuit for the OR-gate can be reused to realize an AND-gate; only the inputs need to be rearranged such that both input gates now receive $U_{\mathrm{in}}^{1}$ and $U_{\mathrm{in}}^{2}$, as shown in Fig. 6(b). The logic for the AND gate is as follows: $\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}\cdot\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}=\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}+\overline{\overline{U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}}}=U_{\mathrm{in}}^{1}\cdot U_{\mathrm{in}}^{2}$. The resulting input-output relation for this AND-gate is shown in Fig. 6(d), confirming that it functions correctly as an AND gate. The dynamic response to slowly varying inputs is comparable to that of the OR gate.
Figure 6: Three combined NAND gates (see text) form (a) an OR gate and (b) an AND-gate. Time traces of the input voltages (blue and orange), representing four consecutive binary states, and the corresponding output voltage (green) for (c) the OR gate and (d) the AND-gate, along with the corresponding truth tables.
To test the repeatability and reproducibility of the gates we need to verify whether they perform as expected over many cycles of the (weakly noisy) input and output sequences. For our continuous system, with output voltages spanning the full range between zero and one volt when the input signal switches, we need to define boundaries separating high- and low-level signals during specific time intervals. Here we set the time interval as the duration between the switching of the input signals and consider an output larger than $0.66\,\mathrm{V}$ as a digital 1 and below $0.33\,\mathrm{V}$ a digital 0. To evaluate the pre-gates, we connect them to the coupling circuit as described above to ensure they function as proper gates. Using these definitions, we tested 100 cycles of all possible input combinations and obtained the following accuracies. Individual XOR-gates operate with accuracy of $\geq 99\%$, NAND-gates with $\geq 99\%$, the combined OR-gate $\geq 99\%$, and the combined AND-gate with $\geq 95\%$ of the time.
VI Conclusion
We designed microfluidic iontronic circuits, mainly based on the combination of volatile memristors and Shinriki-like oscillators, that can perform the non-linear XOR classification task and act as the functional complete NAND logic gate. We showed that these NAND-gates are robust and can be combined to form other gates, especially OR and AND. A key feature of these circuits is their exploitation of noise, here set at the level of $3\,\mathrm{mV}$ on the scale of $1\,\mathrm{V}$ for the signals. This noise level allows for self-excitation of the oscillators from a quiescent (0) to an oscillating (1) state, a feature that stems from the intrinsic non-linearity of the “Memriki” oscillators from which our circuits were composed. We hope that our findings will inspire future explorations of experimental realizations of memristive iontronic devices. While we have not investigated the energy efficiency of our circuits and logic gates in any detail yet, we do not expect their current state to be more energy efficient than present-day silicon technology. In this sense, this work serves as a proof of concept for a potentially versatile iontronic platform for computing, based on the same aqueous electrolyte as mammalian brains.
VII Methods
To evaluate the steady-state conductance of the conical channels, we use the parameters and expressions from [30] for Eq. 1. For a conical channel of length $L=10\,\mathrm{\mu m}$, base radius $R_{b}=200\,\mathrm{nm}$, tip radius $R_{t}=50\,\mathrm{nm}$, and a radius profile $R(x)=R_{t}x/L+R_{b}(1-x/L)$ for $x\in[0,L]$ running from base to tip, the static conductance $g_{\infty}(U)$ is given by
$\frac{g_{\infty}(U)}{g_{0}}$ $=$ (4)
$1+\Delta g\int_{0}^{L}\left(\frac{xR_{t}}{LR(x)}-\frac{\mathrm{e}^{\frac{x\mathrm{Pe}R_{t}^{2}}{LR_{b}R(x)}}-1}{\mathrm{e}^{\mathrm{Pe}\frac{R_{t}}{R_{b}}}-1}\right)\frac{\mathrm{d}x}{L}.$
Here, $g_{0}\approx 4.2\,\mathrm{pS}$ is the zero-field conductance, $\Delta g\approx-3.59$ is the asymmetry parameter, and $\mathrm{Pe}$ is the Peclet number, which varies linearly with the applied voltage $U$ as $\mathrm{Pe}(U)\approx 16U/\mathrm{V}$ for our parameter set (exact expressions provided below). Using the short-hand notation $\Delta R=R_{b}-R_{t}$, and performing the integral, one obtains the explicit static channel conductance
$g_{\infty}(U)=g_{0}\left(1+\Delta g\frac{R_{t}}{\Delta R}F(U)\right)$ (5)
$F(U)=\frac{\mathrm{e}^{\frac{-\mathrm{Pe}R_{t}^{2}}{\Delta RR_{b}}}A(U)}{\mathrm{e}^{\mathrm{Pe}\frac{R_{t}}{R_{b}}}-1}-\left(\frac{R_{b}}{\Delta R}\ln\frac{R_{t}}{R_{b}}+1\right)$ (6)
$A(U)=\mathrm{e}^{\frac{\mathrm{Pe}R_{t}}{\Delta R}}-\frac{R_{b}}{R_{t}}\mathrm{e}^{\frac{\mathrm{Pe}R_{t}^{2}}{\Delta RR_{b}}}+\frac{\Delta R}{R_{t}}+B(U)$ (7)
$B(U)=\frac{-\mathrm{Pe}R_{t}}{\Delta R}\left(\mathrm{Ei}\left(\frac{\mathrm{Pe}R_{t}}{\Delta R}\right)-\mathrm{Ei}\left(\frac{\mathrm{Pe}R_{t}^{2}}{\Delta RR_{b}}\right)\right),$ (8)
where $\mathrm{Ei}(x)=\int_{-\infty}^{x}\mathrm{e}^{-t}/t\mathrm{d}t$ is the exponential integral function.
We stress that this explicit expression for $g_{\infty}(U)$ stems from an (approximate) analytic solution to the fully microscopic Poisson-Nernst-Planck-Stokes equations in the long-channel limit. The dependence of the conductance on $U$ is a direct consequence of electro-osmotic flow, characterized by the dimensionless parameter $\mathrm{Pe}(U)$. This effect relies on the presence of a nonzero surface charge $\sigma$ (and hence a nonzero zeta potential $\psi$ on the channel walls), as becomes explicit by the analytic relation $\mathrm{Pe}(U)=\frac{-eR_{b}}{k_{B}TR_{t}w}U$, where $w=\frac{eD\eta}{k_{B}T\varepsilon\psi}\approx-9.6$ represents the ratio of ionic mobility $D/k_{B}T$ to osmotic mobility $\varepsilon\psi/(e\eta)$. Here $D=1.75\,\mathrm{nm}^{2}/\mathrm{ms}$ is the diffusion coefficient of the monovalent cations and anions in water, $k_{B}$ denotes the Boltzmann constant, $T=293.15\,\mathrm{K}$ is the absolute temperature, and $\varepsilon=80.2\epsilon_{0}\approx 7.1\times 10^{-8}\,\mathrm{F/m}$ is the dielectric constant, and $\eta=1.01\,\mathrm{mPas}$ is the shear viscosity of water. We define the steady zero-field conductance as $g_{0}=\frac{\pi R_{t}R_{b}}{L}\frac{2\rho_{b}e^{2}D}{k_{B}T}$, where $\rho_{b}$ is the ionic bulk concentration, set to $10\,\mathrm{mM}$. The asymmetry parameter is given by $\Delta g=\frac{-2w\Delta R}{R_{b}\mathrm{Du}}\approx-3.59$, where the Dukhin number is $\mathrm{Du}=\sigma/(2\rho_{b}R_{t})\approx-0.25$, with $\sigma=-0.0015\,e/\mathrm{nm}^{2}$ the surface charge density on the channel walls. The corresponding zeta potential is $\psi=\frac{2k_{B}T}{e}\mathrm{asinh}(2\pi\lambda_{D}\lambda_{B}\sigma/e)\approx-10\,\mathrm{mV}$, where $\lambda_{D}=\sqrt{\varepsilon k_{B}T/(2e^{2}\rho_{b})}\approx 30\,\mathrm{nm}$ is the Debye length and $\lambda_{B}=e^{2}/(4\pi\varepsilon k_{B}T)\approx 0.71\,\mathrm{nm}$ is the Bjerrum length.
The characteristic time scale of this conical channel, which is the time scale to build up (or break down) the static voltage-dependent salt concentration profile $\rho_{s}(x;U)$ in the channel, was derived to be the diffusion time $\tau=L^{2}/12D$, valid in the long-channel limit $L\gg R_{b}$. For the present parameters this leads to $\tau\approx 5\,\mathrm{ms}$.
VIII Data availability
All data can be generated with the source-code provided in [38]. The raw data is also available in Ref. [43].
IX Code availability
The code to perform the simulations is provided in [38].
X Competing Interests
There are no competing interests.
XI Author Contributions
N.C.X.S designed the circuits, performed the simulations and analyzed the data. All authors discussed the results and contributed to the manuscript.
XII Acknowledgements
The authors thank Monica Conte and Tim Kamsma for helpful discussions. N.C.X.S and M.D. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. ERC-2019-ADG 884902, SoftML)